OpenStudy (anonymous):

Predicting the kinetic energy of an object from gravitational potential energy?

OpenStudy (anonymous):

A ball that's mass is 5 kg is lifted 1km above the ground, it has GPE of = mgh = 49 kJ of energy. If Work = Fd = 49 kJ F = W/d = 49kJ/ 1000 = 49N of gravitational force.

OpenStudy (anonymous):

acceleration = 9.8 m/s ^2 ( a = F/m)

OpenStudy (anonymous):

Now velocity = d/t

OpenStudy (anonymous):

s = 1/2 at^2 thus, t = 14.2 seconds.

OpenStudy (anonymous):

V = 1000/14.2 = 70.4 m/s

OpenStudy (anonymous):

might want to post this in the science section

OpenStudy (anonymous):

SHOOT!

OpenStudy (anonymous):

LOLZ

OpenStudy (anonymous):

Just realized lol... thanks though.